(a) Given below are the electrical power requirements for five household appliances. Determine the number of kilowatt-hours of electrical energy consumed if all these appliances are running simultaneously for two hours in a house.

solid state colour TV: 145 W

automatic washing machine: 512 W

furnace fan: 500 W

clock: 2 W

humidifier: 177 W

(b) If the average residential cost of electricity is 12 cents per kilowatt-hour, what would be the cost (to the nearest cent) for the energy calculated in (a)?
Answer #1

a) The number of kilowatt-hours of electrical energy consumed if all these appliances are running simultaneously for two hours in a house = (145 + 512 + 500 + 2 + 177)*2 = 1336*2 = 2672 kilowatt-hour

b) the cost (to the nearest cent) for the energy calculated in (a) = 2672*12 cents = 32064 cents